OpenStudy (anonymous):

When using the substitution method, what expression for x in the first equation would be substituted in the second equation? x + 3y = –7 2x + 4y = –4

OpenStudy (anonymous):

x+3y=-7 subtract 3y on both sides x+3y-3y=-7-3y x=-7-3y